OpenStudy (kewlgeek555):

Solve y = -2x + 3 for x.

OpenStudy (amorfide):

Solve y = -2x + 3 for x. you want x on its own first of lets start by getting -2x on its own so you have -2x+3 what do you do to get rid of a +3?

OpenStudy (kewlgeek555):

Um...add? No...subtract or divide. I think divide 2 and 3.

OpenStudy (kewlgeek555):

-2*

OpenStudy (amorfide):

no you have +3 so you want to do the opposite to the sign of the three since you have a positive 3 to get rid of it you want to do the opposite which is subtract 3 but if you do this to the right side you must do it to the left side so re write your equation y-3=-2x+3-3 y-3=-2x now you have -2x you want x on it's own so what do you do to go from -2x to x?

OpenStudy (kewlgeek555):

I divide x from -2?

OpenStudy (amorfide):

no you divide by -2, remember, you want X on its own so you are trying to get rid of the -2 rewrite (y-3)/-2 =-2x/-2 (y-3)/-2=x

OpenStudy (kewlgeek555):

Oh. Okay. Thank you.